# Unified Framework of Forced Magnetic Reconnection and Alfvén Resonance

### v1

@@ V The Ideal Regime
The ideal solution can be recovered from the general solution given by Eq. (28) by taking the limit $Q\gg 1$ (or $\hat{g}\gg S^{-1/3}$) and, simultaneously, $z/Q^{3/2}\ll 1$. The leading order contributions, after imposing the appropriate parity for the stream function, yield
$\tilde{\phi}(\hat{\theta},\hat{g})=\frac{\hat{\psi}_{0}}{2}\frac{e^{-\frac{\hat{g}}{|\hat{k}|}\hat{\theta}}}{\hat{\theta}}.$ (35)
@@
$\Delta_{g}=-\pi\frac{\hat{k}}{\hat{g}}.$
The inverse Fourier transform with respect to $\hat{\theta}$ of Eq. (35) is:
$\hat{\phi}(\hat{x},\hat{g})=i\hat{\psi}_{0}\arctan\left(\frac{|\hat{k}|}{\hat{g}}\hat{x}\right).$ (36)
The solution in Eq. (36) just recovered can be derived also by solving Eq. (20)-(21) by imposing $S=\infty$ (e.g., [6]). The branch cuts and poles of this solution, assuming the forcing given by the standing wave in Eq. (34), is reported in Figure  1. The poles introduced by the boundary condition $\hat{\Xi}(\hat{k},\hat{g})$, in green, provide the long-time finite frequency response which drives the Alfvén resonance. The orange branch cut arises from the complex arctangent contribution and is responsible for the forced reconnection solution when $\hat{\omega}_{0}\rightarrow 0$.
For the standing wave boundary condition, the ideal stream function can be inverted analytically by following the Bromwich contour represented in Figure  1. The solution in terms of the variables $\hat{x}$, $\hat{y}$, and $\hat{t}$ is given by
$\phi=-\Phi\left\{\cos(\hat{\omega}_{0}\hat{t})\left[CI(z_{+})-CI(z_{-})+\ln\left(\frac{z_{-}}{z_{+}}\right)\right]-\frac{2\sin(\hat{k}_{0}\hat{x}\hat{t})}{\hat{\omega}_{0}\hat{t}}+\sin(\hat{\omega}_{0}\hat{t})\left[SI(z_{+})+SI(z_{-})\right]\right\},$ (37)
@@
$z_{\pm}=(\hat{k}_{0}\hat{x}\pm\hat{\omega}_{0})\hat{t},\quad\Phi=\frac{\Xi_{0}}{\pi}\frac{\hat{\omega}_{0}\sin[\hat{k}_{0}\hat{y}]}{\sinh[\hat{k}_{0}]},$ (38)
and $SI(z_{\pm})$, $CI(z_{\pm})$ are the sine and cosine integral functions, respectively. We see that for $\hat{\omega}_{0}\neq 0$ we recover the expected logarithmic singularity at the resonant site, where the phase velocity of the driven boundary perturbation, $v_{ph}=\hat{\omega}_{0}/\hat{k}_{0}$, is equal to the local Alfvén speed. The logarithmic contribution to the stream function originates from the residues at the poles $\hat{g}=\pm\hat{\omega}_{0}$, while the sine and cosine integrals arise from evaluating the jump of the Laplace integral across the branch cut. For long times, the contribution of the logarithmic response to the resonant poles dominates over that of the sine and cosine integrals, which initially provide much of the oscillatory behavior of the solution. In the limit $\hat{\omega}_{0}\rightarrow 0$, Eq. (37) describes forced reconnection at the neutral line and recovers the results of [2].
@@
$\tilde{J}_{z}=\frac{\hat{\psi}_{0}}{2}\frac{\hat{k}}{\hat{g}}\left(\frac{\hat{g}}{|\hat{k}|}\hat{\theta}+1\right)e^{-\frac{\hat{g}}{|\hat{k}|}\hat{\theta}},$ (39)
@@
$J_{z}=-\Xi_{0}\frac{2}{\pi}\frac{\hat{k}_{0}^{4}\cos(\hat{k}_{0}\hat{y})}{\sinh(\hat{k}_{0})}\frac{\hat{x}^{2}\hat{t}}{(\hat{k}_{0}^{2}\hat{x}^{2}-\hat{\omega}_{0}^{2})}\left[\cos(\hat{k}_{0}\hat{x}\hat{t})+\frac{(\hat{k}_{0}^{2}\hat{x}^{2}-3\hat{\omega}_{0}^{2})}{(\hat{k}_{0}^{2}\hat{x}^{2}-\hat{\omega}_{0}^{2})}\frac{\sin(\hat{k}_{0}\hat{x}\hat{t})}{\hat{k}_{0}\hat{x}\hat{t}}+\frac{2\hat{\omega}_{0}^{3}}{\hat{k}_{0}^{2}\hat{x}^{2}\hat{t}}\frac{\sin(\hat{\omega}_{0}\hat{t})}{(\hat{k}_{0}^{2}\hat{x}^{2}-\hat{\omega}_{0}^{2})}\right].$ (40)
For $\hat{t}\ll 1$ Eq. (40) is given by
$J_{z}=-\Xi_{0}\frac{4}{\pi}\frac{\hat{k}_{0}^{2}\cos(\hat{k}_{0}\hat{y})}{\sinh(\hat{k}_{0})}\hat{t}\left[1+\frac{1}{6}(\hat{k}_{0}^{2}\hat{x}^{2}+\hat{\omega}_{0}^{2})\hat{t}^{2}\right]+\mathcal{O}(\hat{t}^{4}).$ (41)
As can be seen from Eq. (41), the current density’s amplitude increases linearly with time regardless of the forcing frequency to lowest order. As time goes by, the more general Eq. (40) recovers the result of [2] for $\hat{\omega}_{0}=0$, where the amplitude of $J_{z}$ increases linearly with time at $\hat{x}=0$, and it is localized in a region whose width decreases as $1/\hat{t}$. Outside the origin, the current density’s amplitude still increases linearly with time but its spatial behavior is highly oscillatory. For $\hat{\omega}_{0}\neq 0$, the current density at the origin is a purely oscillating function. For long times, however, it grows faster than linear near the resonant point where it builds up by eventually leading to a singularity.
@@ VI The Resistive Regime
We separate now the discussion of the time-asymptotic behavior, corresponding to $Q\ll 1$, into two cases: the case of small or zero frequency, $\hat{\omega}_{0}\ll S^{-1/3}$, for which we expect to describe forced magnetic reconnection, and the case of high frequency $\hat{\omega}_{0}\gg S^{-1/3}$.
@@ VI.1 Forced Reconnection
In the limit $Q\ll 1$ (or $\hat{g}\ll S^{-1/3}\hat{k}^{2/3}$) Eq. (28) with Eq. (30) approximates to
$\tilde{\phi}=-\frac{\Gamma[\frac{3}{4}]}{2\pi}\hat{\psi}_{0}\frac{\hat{g}}{\hat{k}}\left(\frac{1}{|\hat{k}|}\sqrt{\frac{\hat{g}}{S}}\hat{\theta}^{2}\right)^{1/4}K_{\frac{1}{4}}\left(\frac{1}{|\hat{k}|}\sqrt{\frac{\hat{g}}{S}}\frac{\hat{\theta}^{2}}{2}\right),$ (42)
@@
$\Delta_{g}=\sqrt{2}(\frac{\hat{g}^{5}S^{3}}{\hat{k}^{2}})^{1/4}\Gamma(3/4)^{2}.$ (43)
From the definition of $\Delta_{g}$, we observe the time scale $\hat{t}\sim S^{3/5}$ after which reconnection reaches a steady state [2, 3]. The flux function at the origin $\hat{\psi}_{0}$, given below, is consistent with that found by [2, 3], albeit for different boundary conditions,
$\hat{\psi}_{0}$ $=$ $\frac{\frac{2\hat{k}}{\sinh(\hat{k})}\hat{\Xi}}{\sqrt{2}(\frac{\hat{g}^{5}S^{3}}{\hat{k}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{\hat{k}}{\tanh(\hat{k})}}.$ (44)
Figure 2: The branches and poles of the constant-psi solution (45) are shown here in the complex plane. In green are the resonant poles of the boundary condition, in blue the singular contributions of the $\hat{g}^{5/4}$ root, and in red is the branch of the Modified Bessel and Struve functions which extends from $\hat{g}=-\infty$ to $\hat{g}=0$.
@@
$\hat{\phi}=\Xi_{0}\sqrt{\frac{\pi}{2}}\Gamma(\frac{3}{4})\frac{\sin(\hat{k}_{0}\hat{y})}{\sinh(\hat{k}_{0})}\frac{\hat{g}^{2}}{(\hat{g}^{2}+\hat{\omega}_{0}^{2})}\left(\hat{k}_{0}\sqrt{\frac{S}{\hat{g}}}\right)^{3/4}\sqrt{\hat{x}}\left(\frac{I_{1/4}\left(\hat{k}_{0}\sqrt{\frac{S}{\hat{g}}}\frac{\hat{x}^{2}}{2}\right)-L_{1/4}\left(\hat{k}_{0}\sqrt{\frac{S}{\hat{g}}}\frac{\hat{x}^{2}}{2}\right)}{\sqrt{2}(\frac{\hat{g}^{5}S^{3}}{\hat{k}_{0}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{\hat{k}_{0}}{\tanh(\hat{k}_{0})}}\right),$ (45)
@@
where $I_{\nu}$ is the modified Bessel function of the first kind and $L_{\nu}$ is the modified Struve function. The branch and poles of this solution with the Bromwich contour are sketched in Figure 2. The green poles represent once more the resonant contributions of the forcing that merge with the branch cut, represented in red, in the limit of $\omega_{0}\rightarrow 0$. The blue poles are the roots of the denominator in the last term in Eq. (45). It is noted that we recover the same branches and poles as [2], but in their case the resonant poles of the boundary overlap with the branch cut at $\hat{g}=0$.
@@
To find the time dependent stream function in this regime, we use Mathematica to numerically take the inverse Laplace transform of Eq. (45) about the Bromwich contour just discussed. In the limit $\omega_{0}\rightarrow 0$, the contributions to the integral are the residues of Eq. (45) evaluated at the $\hat{g}^{5/4}$ poles (the blue poles in Fig. 2) and the jump across the branch extending from $\hat{g}=-\infty$ to $\hat{g}=0$. For small but non zero frequencies $\hat{\omega}_{0}$, we must include also the contribution of the residues at $\hat{g}=\pm i\hat{\omega}_{0}$, which add an oscillatory contribution to the solution that is non-negligible for $\hat{t}\rightarrow\infty$.
@@
The resulting stream function $\phi(\hat{x},\hat{y},\hat{t})$ is compared with results from a numerical simulation in Figure 3, where dashed lines correspond to the theoretical solutions and solid lines correspond to results from ab-initio numerical simulations. Details on the numerical code can be found in the Appendix. The constant-psi solution is represented by the purple, cyan and green colors for $\hat{\omega}_{0}=0$, $\hat{\omega}_{0}=0.0002$, and $\hat{\omega}_{0}=0.002$, respectively, where we have fixed $\hat{y}=\pi/2$ and $\hat{t}=3000$, which is sufficiently long compared to the resistive time scale $\hat{t}\sim S^{1/3}$. The chosen frequencies all satisfy the condition $\hat{\omega}_{0}\ll S^{-1/3}$ and, as can be seen, the solution corresponding to the constant-psi regime matches very well with the simulation results for each frequency.
@@
Figure 3: The time-asymptotic solution for the stream function $\phi(\hat{x},\hat{y},\hat{t})$ (dashed lines) compared with numerical simulations (solid lines). The dashed green, cyan, and purple lines are the inverse Laplace transformed constant-psi solutions, Eq. (45), for $\omega_{0}=0.002$, $\omega_{0}=0.0002$, and $\omega_{0}=0$, respectively. These solutions are best seen in the inset plot, showing an enlarged picture of the region around the neutral line. The dashed blue and red lines are the inverse Laplace transformed Alfvén resonance solutions, Eq. (51), for $\omega_{0}=0.1$ and $\omega_{0}=0.05$, respectively. The constant-psi solution for $\omega_{0}=0.1$ has been also plotted (dotted line) to demonstrate the breakdown of the constant-psi solution as the system transitions towards Alfvén resonance for increasing frequencies.
@@
The current density is found by taking the inverse Fourier transform of (22) in the limit $Q\ll\bar{\theta}^{2}$
$\hat{J}_{z}=-(S\hat{k})\int^{\infty}_{-\infty}e^{i\hat{x}\hat{\theta}}\frac{d}{d\hat{\theta}}\tilde{\phi},$ (46)
@@
$\hat{J}_{z}=-i(S\hat{k})\hat{x}\hat{\phi},$ (47)
@@
$\hat{J}_{z}=\Xi_{0}\sqrt{\frac{\pi}{2}}\Gamma(\frac{3}{4})\frac{\hat{k}_{0}\cos(\hat{k}_{0}\hat{y})}{\sinh(\hat{k}_{0})}\frac{\hat{g}^{2}S}{(\hat{g}^{2}+\hat{\omega}_{0}^{2})}\left(\hat{k}_{0}\sqrt{\frac{S}{\hat{g}}}\hat{x}^{2}\right)^{3/4}\left(\frac{I_{1/4}\left(\hat{k}_{0}\sqrt{\frac{S}{\hat{g}}}\frac{\hat{x}^{2}}{2}\right)-L_{1/4}\left(\hat{k}_{0}\sqrt{\frac{S}{\hat{g}}}\frac{\hat{x}^{2}}{2}\right)}{\sqrt{2}(\frac{\hat{g}^{5}S^{3}}{\hat{k}_{0}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{\hat{k}_{0}}{\tanh(\hat{k}_{0})}}\right).$ (48)
[…]
@@
In principle, Eq. (48) describes also the current density at the origin under an appropriate limit by expanding the Modified Bessel and Struve functions, $I_{1/4}$ and $L_{1/4}$ respectively, for small arguments. For simplicity, to find the current density at $\hat{x}=0$ required to calculate the reconnection rate, we follow the method adopted in other works by employing the relation
$J_{z}(0,\hat{y},\hat{t})=S\frac{\partial\psi_{0}}{\partial\hat{t}}=S\mathcal{L}^{-1}(\hat{g}\hat{\psi}_{0}),$ (49)
@@
obtainable from (46) by imposing $\hat{x}=0$ in the exponential, making use of the relationship of $\tilde{\phi}$ and $\tilde{\psi}$, Eq. (19), and, finally, by applying the limit $Q\ll\bar{\theta}^{2}$. We have also assumed here that reconnection has been developed sufficiently such that $\psi$ is approximately constant across the layer, as is the case for the constant-psi approximation. For the boundary condition employed in this paper, Eq. (49) yields,
@@
$\hat{J}_{z}(0,\hat{y},\hat{g})=-\Xi_{0}\frac{2\hat{k}_{0}\cos(\hat{k}_{0}\hat{y})}{\sinh(\hat{k}_{0})}\frac{\hat{g}^{2}S}{(\hat{g}^{2}+\hat{\omega}_{0}^{2})}\left(\frac{1}{\sqrt{2}(\frac{\hat{g}^{5}S^{3}}{\hat{k}_{0}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{\hat{k}_{0}}{\tanh(\hat{k}_{0})}}\right).$ (50)
@@
We now evaluate numerically its inverse Laplace transform via the Bromwich contour in Figure 2 (using Mathematica) to calculate the reconnection rate $R(\hat{t})=J_{z}(0,\hat{t})/S$ at a given position $\hat{y}$.
@@
The time evolution of the reconnection rate $R(\hat{t})$ in the cases of $\hat{\omega}_{0}=0$ and $\hat{\omega}_{0}=0.002$ is plotted in Figure 4, left and middle panels, where we compare $R$ from the 2D simulation (orange dashed line) with the ideal (blue) and constant-psi (green) solutions. The reconnection rate behaves as expected for the zero frequency case, (left panel): it initially grows linearly with time as described by the ideal solution and for $\hat{t}\gg S^{1/3}$ it follows the constant-psi solution. After times much longer than the reconnection time scale $\hat{t}\sim S^{3/5}$, the reconnection rate approaches zero as the flux function at the origin reaches a steady state. The reconnection rate has a first peak in the intermediate regime, between the ideal and constant-psi regimes, when the resistive layer forms at $\hat{t}\sim S^{1/3}$. In the case of a small but non zero frequency (middle panel), the reconnection rate follows an analogous evolution but, over a longer time scale, instead of reaching a steady state the reconnection rate oscillates in response to the boundary driver. The fact that the small but non zero frequency case also gives rise to reconnection isn’t surprising, as discussed in the Introduction, and was analyzed previously by [25] for the eigenvalue problem and by [26] in the study of the interaction of resonant magnetic perturbations with rotating plasmas. In a later section we discuss in more details the effects of finite frequency on reconnection.
@@
Figure 4: The left and middle panels show the reconnection rate as a function of time from numerical simulations compared with the two asymptotic solutions: the ideal solution, Eq. (40) divided by $S$ in blue, and the inverse Laplace transform of the constant-psi solution, Eq. (49) in green. We show results for $\hat{\omega}_{0}=0$ (left panel) and $\hat{\omega}_{0}=0.002$ (middle panel). In the right panel, the simulated current density for $\hat{\omega}_{0}=0.1$ has been plotted at the resonant site. It is compared with the same ideal solution (blue color) and with the resistive solution solution given by the inverse Fourier transform of Eq. (54) (green color). The parameters used here are $S=10^{6}$, $\hat{k}_{0}=1$, $\hat{y}=\pi/2$.

### v2

@@ V The Ideal Regime
The ideal solution can be recovered from the general solution given by Eq. (28) by taking the limit $Q\gg 1$ (or $g\gg S^{-1/3}$) and, simultaneously, $z/Q^{3/2}\ll 1$. The leading order contributions, after imposing the appropriate parity for the stream function, yield
$\tilde{\phi}(\theta,k,g)=\frac{\hat{\psi}_{0}}{2}\frac{e^{-\frac{g}{|k|}\theta}}{\theta}.$ (35)
@@
$\Delta_{g}=-\pi\frac{k}{g}.$
The inverse Fourier transform with respect to $\theta$ of Eq. (35) is:
$\hat{\phi}(x,k,g)=i\hat{\psi}_{0}\arctan\left(\frac{|k|}{g}x\right).$ (36)
The solution in Eq. (36) just recovered can be derived also by solving Eq. (20)-(21) by imposing $S=\infty$ (e.g., [6]). The branch cuts and poles of this solution, assuming the forcing given by the standing wave in Eq. (34), is reported in Figure  1. The poles introduced by the boundary condition $\hat{\Xi}(k,g)$, in green, provide the long-time finite frequency response which drives the Alfvén resonance. The orange branch cut arises from the complex arctangent contribution and is responsible for the forced reconnection solution when $\omega_{0}\rightarrow 0$.
For the standing wave boundary condition, the ideal stream function can be inverted analytically by following the Bromwich contour represented in Figure  1. The solution in terms of the variables $x$, $y$, and $t$ is given by
$\phi=-\Phi\left\{\cos(\omega_{0}t)\left[CI(z_{+})-CI(z_{-})+\ln\left(\frac{z_{-}}{z_{+}}\right)\right]-\frac{2\sin(k_{0}xt)}{\omega_{0}t}+\sin(\omega_{0}t)\left[SI(z_{+})+SI(z_{-})\right]\right\},$ (37)
@@
$z_{\pm}=(k_{0}x\pm\omega_{0})t,\quad\Phi=\frac{\Xi_{0}}{\pi}\frac{\omega_{0}\sin[k_{0}y]}{\sinh[k_{0}]},$ (38)
and $SI(z_{\pm})$, $CI(z_{\pm})$ are the sine and cosine integral functions, respectively. We see that for $\omega_{0}\neq 0$ we recover the expected logarithmic singularity at the resonant site, where the phase velocity of the driven boundary perturbation, $v_{ph}=\omega_{0}/k_{0}$, is equal to the local Alfvén speed. The logarithmic contribution to the stream function originates from the residues at the poles $g=\pm\omega_{0}$, while the sine and cosine integrals arise from evaluating the jump of the Laplace integral across the branch cut. For long times, the contribution of the logarithmic response to the resonant poles dominates over that of the sine and cosine integrals, which initially provide much of the oscillatory behavior of the solution. In the limit $\omega_{0}\rightarrow 0$, Eq. (37) describes forced reconnection at the neutral line and recovers the results of [2].
@@
$\tilde{J}_{z}=\frac{\psi_{0}}{2}\frac{k}{g}\left(\frac{g}{|k|}\theta+1\right)e^{-\frac{g}{|k|}\theta},$ (39)
@@
$J_{z}=-\Xi_{0}\frac{2}{\pi}\frac{k_{0}^{4}\cos(k_{0}y)}{\sinh(k_{0})}\frac{x^{2}t}{(k_{0}^{2}x^{2}-\omega_{0}^{2})}\left[\cos(k_{0}xt)+\frac{(k_{0}^{2}x^{2}-3\omega_{0}^{2})}{(k_{0}^{2}x^{2}-\omega_{0}^{2})}\frac{\sin(k_{0}xt)}{k_{0}xt}+\frac{2\omega_{0}^{3}}{k_{0}^{2}x^{2}t}\frac{\sin(\omega_{0}t)}{(k_{0}^{2}x^{2}-\omega_{0}^{2})}\right].$ (40)
For $t\ll 1$ Eq. (40) is given by
$J_{z}=-\Xi_{0}\frac{4}{\pi}\frac{k_{0}^{2}\cos(k_{0}y)}{\sinh(k_{0})}t\left[1+\frac{1}{6}(k_{0}^{2}x^{2}+\omega_{0}^{2})t^{2}\right]+\mathcal{O}(t^{4}).$ (41)
As can be seen from Eq. (41), the current density’s amplitude increases linearly with time regardless of the forcing frequency to lowest order. As time goes by, the more general Eq. (40) recovers the result of [2] for $\omega_{0}=0$, where the amplitude of $J_{z}$ increases linearly with time at $x=0$, and it is localized in a region whose width decreases as $1/t$. Outside the origin, the current density’s amplitude still increases linearly with time but its spatial behavior is highly oscillatory. For $\omega_{0}\neq 0$, the current density at the origin is a purely oscillating function. For long times, however, it grows faster than linear near the resonant point where it builds up by eventually leading to a singularity.
@@ VI The Resistive Regime
We separate now the discussion of the time-asymptotic behavior, corresponding to $Q\ll 1$, into two cases: the case of small or zero frequency, $\omega_{0}\ll S^{-1/3}$, for which we expect to describe forced magnetic reconnection, and the case of high frequency $\omega_{0}\gg S^{-1/3}$.
@@ VI.1 Forced Reconnection
In the limit $Q\ll 1$ (or $g\ll S^{-1/3}k^{2/3}$) Eq. (28) with Eq. (30) approximates to
$\tilde{\phi}=-\frac{\Gamma[\frac{3}{4}]}{2\pi}\psi_{0}\frac{g}{k}\left(\frac{1}{|k|}\sqrt{\frac{g}{S}}\theta^{2}\right)^{1/4}K_{\frac{1}{4}}\left(\frac{1}{|k|}\sqrt{\frac{g}{S}}\frac{\theta^{2}}{2}\right),$ (42)
@@
$\Delta_{g}=\sqrt{2}(\frac{g^{5}S^{3}}{k^{2}})^{1/4}\Gamma(3/4)^{2}.$ (43)
From the definition of $\Delta_{g}$, we observe the time scale $t\sim S^{3/5}$ after which reconnection reaches a steady state [2, 3]. The flux function at the origin $\hat{\psi}_{0}$, given below, is consistent with that found by [2, 3], albeit for different boundary conditions,
$\hat{\psi}_{0}$ $=$ $\frac{\frac{2k}{\sinh(k)}\hat{\Xi}}{\sqrt{2}(\frac{g^{5}S^{3}}{k^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{k}{\tanh(k)}}.$ (44)
Figure 2: The branches and poles of the constant-psi solution (45) are shown here in the complex plane. In green are the resonant poles of the boundary condition, in blue the singular contributions of the $g^{5/4}$ root, and in red is the branch of the Modified Bessel and Struve functions which extends from $g=-\infty$ to $g=0$.
@@
$\hat{\phi}=\Xi_{0}\sqrt{\frac{\pi}{2}}\Gamma(\frac{3}{4})\frac{\sin(k_{0}y)}{\sinh(k_{0})}\frac{g^{2}}{(g^{2}+\omega_{0}^{2})}\left(k_{0}\sqrt{\frac{S}{g}}\right)^{3/4}\sqrt{x}\left(\frac{I_{1/4}\left(k_{0}\sqrt{\frac{S}{g}}\frac{x^{2}}{2}\right)-L_{1/4}\left(k_{0}\sqrt{\frac{S}{g}}\frac{x^{2}}{2}\right)}{\sqrt{2}(\frac{g^{5}S^{3}}{k_{0}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{k_{0}}{\tanh(k_{0})}}\right),$ (45)
@@
where $I_{\nu}$ is the modified Bessel function of the first kind and $L_{\nu}$ is the modified Struve function. The branch and poles of this solution with the Bromwich contour are sketched in Figure 2. The green poles represent once more the resonant contributions of the forcing that merge with the branch cut, represented in red, in the limit of $\omega_{0}\rightarrow 0$. The blue poles are the roots of the denominator in the last term in Eq. (45). It is noted that we recover the same branches and poles as [2], but in their case the resonant poles of the boundary overlap with the branch cut at $g=0$.
@@
To find the time dependent stream function in this regime, we use Mathematica to numerically take the inverse Laplace transform of Eq. (45) about the Bromwich contour just discussed. In the limit $\omega_{0}\rightarrow 0$, the contributions to the integral are the residues of Eq. (45) evaluated at the $g^{5/4}$ poles (the blue poles in Fig. 2) and the jump across the branch extending from $g=-\infty$ to $g=0$. For small but non zero frequencies $\omega_{0}$, we must include also the contribution of the residues at $g=\pm i\omega_{0}$, which add an oscillatory contribution to the solution that is non-negligible for $t\rightarrow\infty$.
@@
The resulting stream function $\phi(x,y,t)$ is compared with results from a numerical simulation in Figure 3, where dashed lines correspond to the theoretical solutions and solid lines correspond to results from ab-initio numerical simulations. Details on the numerical code can be found in the Appendix. The constant-psi solution is represented by the purple, cyan and green colors for $\omega_{0}=0$, $\omega_{0}=0.0002$, and $\omega_{0}=0.002$, respectively, where we have fixed $y=\pi/2$ and $t=3000$, which is sufficiently long compared to the resistive time scale $t\sim S^{1/3}$. The chosen frequencies all satisfy the condition $\omega_{0}\ll S^{-1/3}$ and, as can be seen, the solution corresponding to the constant-psi regime matches very well with the simulation results for each frequency.
@@
Figure 3: The time-asymptotic solution for the stream function $\phi(x,y,t)$ (dashed lines) compared with numerical simulations (solid lines). The dashed green, cyan, and purple lines are the inverse Laplace transformed constant-psi solutions, Eq. (45), for $\omega_{0}=0.002$, $\omega_{0}=0.0002$, and $\omega_{0}=0$, respectively. These solutions are best seen in the inset plot, showing an enlarged picture of the region around the neutral line. The dashed blue and red lines are the inverse Laplace transformed Alfvén resonance solutions, Eq. (51), for $\omega_{0}=0.1$ and $\omega_{0}=0.05$, respectively. The constant-psi solution for $\omega_{0}=0.1$ has been also plotted (dotted line) to demonstrate the breakdown of the constant-psi solution as the system transitions towards Alfvén resonance for increasing frequencies.
@@
The current density is found by taking the inverse Fourier transform of (22) in the limit $Q\ll\theta^{2}$
$\hat{J}_{z}=-(Sk)\int^{\infty}_{-\infty}e^{ix\theta}\frac{d}{d\theta}\tilde{\phi},$ (46)
@@
$\hat{J}_{z}=-i(Sk)x\hat{\phi}(x,k,g),$ (47)
from the definition of the Fourier transform. By using the appropriate $\hat{\phi}$ for this limit for constant-psi (from equation (45) in Fourier $k$ space) $\hat{J_{z}}$ is given by
$\hat{J}_{z}=\Xi_{0}\sqrt{\frac{\pi}{2}}\Gamma(\frac{3}{4})\frac{k_{0}\cos(k_{0}y)}{\sinh(k_{0})}\frac{g^{2}S}{(g^{2}+\omega_{0}^{2})}\left(k_{0}\sqrt{\frac{S}{g}}x^{2}\right)^{3/4}\left(\frac{I_{1/4}\left(k_{0}\sqrt{\frac{S}{g}}\frac{x^{2}}{2}\right)-L_{1/4}\left(k_{0}\sqrt{\frac{S}{g}}\frac{x^{2}}{2}\right)}{\sqrt{2}(\frac{g^{5}S^{3}}{k_{0}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{k_{0}}{\tanh(k_{0})}}\right).$ (48)
@@
In principle, Eq. (48) describes also the current density at the origin under an appropriate limit by expanding the Modified Bessel and Struve functions, $I_{1/4}$ and $L_{1/4}$ respectively, for small arguments. For simplicity, to find the current density at $x=0$ required to calculate the reconnection rate, we follow the method adopted in other works by employing the relation
$J_{z}(0,y,t)=S\frac{\partial\psi_{0}}{\partial t}=S\mathcal{L}^{-1}(g\hat{\psi}_{0}),$ (49)
@@
obtainable from (46) by imposing $x=0$ in the exponential, making use of the relationship of $\tilde{\phi}$ and $\tilde{\psi}$, Eq. (19), and, finally, by applying the limit $Q\ll\theta^{2}$. We have also assumed here that reconnection has been developed sufficiently such that $\psi$ is approximately constant across the layer, as is the case for the constant-psi approximation. For the boundary condition employed in this paper, Eq. (49) yields,
@@
$\hat{J}_{z}(0,y,g)=-\Xi_{0}\frac{2k_{0}\cos(k_{0}y)}{\sinh(k_{0})}\frac{g^{2}S}{(g^{2}+\omega_{0}^{2})}\left(\frac{1}{\sqrt{2}(\frac{g^{5}S^{3}}{k_{0}^{2}})^{1/4}\Gamma(3/4)^{2}+2\frac{k_{0}}{\tanh(k_{0})}}\right).$ (50)
@@
We now evaluate numerically its inverse Laplace transform via the Bromwich contour in Figure 2 (using Mathematica) to calculate the reconnection rate $R(t)=J_{z}(0,t)/S$ at a given position $y$.
@@
The time evolution of the reconnection rate $R(t)$ in the cases of $\omega_{0}=0$ and $\omega_{0}=0.002$ is plotted in Figure 4, left and middle panels, where we compare $R$ from the 2D simulation (orange dashed line) with the ideal (blue) and constant-psi (green) solutions. The reconnection rate behaves as expected for the zero frequency case, (left panel): it initially grows linearly with time as described by the ideal solution and for $t\gg S^{1/3}$ it follows the constant-psi solution. After times much longer than the reconnection time scale $t\sim S^{3/5}$, the reconnection rate approaches zero as the flux function at the origin reaches a steady state. The reconnection rate has a first peak in the intermediate regime, between the ideal and constant-psi regimes, when the resistive layer forms at $t\sim S^{1/3}$. In the case of a small but non zero frequency (middle panel), the reconnection rate follows an analogous evolution but, over a longer time scale, instead of reaching a steady state the reconnection rate oscillates in response to the boundary driver. The fact that the small but non zero frequency case also gives rise to reconnection isn’t surprising, as discussed in the Introduction, and was analyzed previously by [25] for the eigenvalue problem and by [26] in the study of the interaction of resonant magnetic perturbations with rotating plasmas. In a later section we discuss in more details the effects of finite frequency on reconnection.
@@
Figure 4: The left and middle panels show the reconnection rate as a function of time from numerical simulations compared with the two asymptotic solutions: the ideal solution, Eq. (40) divided by $S$ in blue, and the inverse Laplace transform of the constant-psi solution, Eq. (49) in green. We show results for $\omega_{0}=0$ (left panel) and $\omega_{0}=0.002$ (middle panel). In the right panel, the simulated current density for $\omega_{0}=0.1$ has been plotted at the resonant site. It is compared with the same ideal solution (blue color) and with the resistive solution solution given by the inverse Fourier transform of Eq. (54) (green color). The parameters used here are $S=10^{6}$, $k_{0}=1$, $y=\pi/2$.